OpenStudy (anonymous):

How many solutions does this linear system have? y = -x + 2 x + y =5

OpenStudy (bahrom7893):

y = -x+2 y = -x +5 again parallel lines, no solutions

OpenStudy (anonymous):

Thanks :-)